Betty walks 2km due west then 2km due south calculate the displacement


1
Expert's answer
2021-10-31T18:15:51-0400

"s_w= 2km"

"s_s= 2km"

"\vec{W}\perp \vec S"

"s= \sqrt{s_w^2+s_s^2}=\sqrt{2^2+2^2}\approx2.83km"


"\text{Answer: }2.83km"